OpenStudy (anonymous):

solve for the higher order derivative. 3 find "y" if y = __________ (x +1)^4 please help me to solve this question.

OpenStudy (anonymous):



OpenStudy (anonymous):

write it as \[3(x+1)^{-4}\] and use the all mighty power rule

OpenStudy (anonymous):

got that? should take one step only

OpenStudy (anonymous):

@satellite73 how?

OpenStudy (anonymous):

\[\frac{d}{dx}x^n=nx^{n-1}\] in your case \(n=-4\)

mathslover (mathslover):

I'm not sure whether user is acquainted with the formulas for finding derivatives and all. In the previous question, asked by the same user, I found the asker not getting the formulas and all. Sam, before Satellite goes on, I just want to tell us about what you've read till now. Thanks!

OpenStudy (anonymous):

@SamMarie14 do you know the power rule?

OpenStudy (anonymous):

ower Rule xn nxn−1

OpenStudy (anonymous):

ok good use it in this case where you have \(3(x+1)^{-4}\)

OpenStudy (anonymous):

@satellite73 how sir?

OpenStudy (anonymous):

drop the \(-4\) out front as a multiplier, then subtract one from it

OpenStudy (anonymous):

\[\huge 3(x+1)^{\color{red}{-4}}\] \[\huge \color{red}{-4}\times 3(x+1)^{\color{red}{-4}-1}\]

OpenStudy (anonymous):

-12(x + 1)^-3 @satellite73

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

what is \[\huge -4-1\]?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

are you taking a calculus class ?

OpenStudy (anonymous):

i forgot the sign -5

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

now that you have \[y'=-12(x+1)^{-5}\] if you want \[y''\] you have to do it again

OpenStudy (freckles):

when you say "y" this each time you do mean just y" right y" means second derivative

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

can you help me again sir @freckles ?

OpenStudy (anonymous):

@satellite73 how can i get that y?

OpenStudy (anonymous):

same thing again repeat the process as before

OpenStudy (jhannybean):

Use the power rule again.

OpenStudy (anonymous):

-5 x -12 (x+1)^-5 -1? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-60(x+1)^-6

OpenStudy (anonymous):

watch your signs

OpenStudy (anonymous):

ogh sorry

OpenStudy (jhannybean):

And so \(y'' = 60(x+1)^{-6} = \dfrac{60}{(x+1)^6} \)

OpenStudy (anonymous):

@satellite73 60(x+1)^-6 then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Jhannybean thats it?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Yawp.

OpenStudy (anonymous):

@Jhannybean thats the final?